OpenStudy (anonymous):

CAN SOMEONE CHECK MY ANSWER? Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, five divided by three, negative five divided by eighteen, ... I got: Diverges Here are my choices: Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (anonymous):

hard to read but it looks like they go to zero

OpenStudy (anonymous):

\[60,-10,\frac{5}{3},-\frac{5}{18}...\]right?

OpenStudy (anonymous):

is it clear what you are doing to one number in the sequence to get the next number?

OpenStudy (anonymous):

Correct @satellite73

OpenStudy (anonymous):

Dividing by -6

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and if you keep dividing by -6, what happens ?

OpenStudy (anonymous):

The numbers get smaller

OpenStudy (anonymous):

Eventually to 0?

OpenStudy (anonymous):

like for example \[\frac{5}{(-6)^{20}}\]

OpenStudy (anonymous):

yeah goes to zero (lickety split i might add)

OpenStudy (anonymous):

Oh, wow you made it seem very simple. Thank you very much for your time!

OpenStudy (anonymous):

lol you are welcome, it is rather simple, once you know what you are doing

OpenStudy (anonymous):

unlike for example the question @Loser66 answered earlier

OpenStudy (anonymous):

which was very simple once you knew the answer, but not so simple beforehand